a 26-inch pieve of steel is cut into three pieces so that the second piece is twice as long the first piece, and the third piece is two inches more than three times the length of the first piece. Find the lengths of the pieces.
Answer by mananth(16949)

a 26-inch pieve of steel is cut into three pieces so that the second piece is twice as long the first piece, and the third piece is two inches more than three times the length of the first piece. Find the lengths of the pieces.
Let the length of first piece be assumed as x
the second piece is twice the first so it is 2x
The third piece is 3x+2
The sum of all the theree pieces is 26 inches
x+2x+3x+2=26
6x= 26-2
x=24/6
x=4
second poece will be 4*2=8 inches
The third piece will be 3x+2 = 3*4+2
=14 inches
4+8+14= 26 inches